OpenStudy (anonymous):

Find the equestion of the line that passes through the point (4,-1) and (5,-3)

OpenStudy (anonymous):

slope = (-3-(-1))/(5-4)=-2 y=-2x+a -3=-10+a a=7 y=-2x+7

OpenStudy (anonymous):

slope = sub of y / sub of x then plug in the point in y=mx+a

OpenStudy (anonymous):

sorry how do you plug in the y=mx+a? thats not the answer? sorry I feel stupid I am just confuzed

OpenStudy (anonymous):

What form you write the equation in matters. Is it "slope intercept" form?

OpenStudy (anonymous):

Slope = (y2 - y1)/(x2 - x1) = (-3+1)(5-4) = -2 y = -2x + b Now, passing through 4, -1, you have b = -1 + 2(4) = 7 So, the equation is: y = -2x + 7

OpenStudy (anonymous):

You can also re-write it as: 2x + y = 7